OpenStudy (anonymous):

evauluate the expression e^(ln18)

OpenStudy (anonymous):

e^ln a = a

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

Just a simple evaluation

OpenStudy (anonymous):

e^x is the inverse function of ln x so e^ln x = x

OpenStudy (anonymous):

Example e^ln 10 = 10